OpenStudy (anonymous):

Find all roots and check your answer 5x^3-23x=57x

OpenStudy (anonymous):

5x^3 - 23x - 57x = 0 5x^3 - 80x = 0 x(5x^2 - 80) = 0 so x = 0 is one solution 5x^2 - 80 = 0 x^2 - 16 = 0 x = sqrt(16) so x = -4, +4 is two other solutions

OpenStudy (anonymous):

can i ask you another one

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

2(x+15)=9x+2(x^2-15)

OpenStudy (anonymous):

2x + 30 = 2x^2 - 30 + 9x moving to the right 2x^2 + 7x - 60 = 0 (2x + 15)(x - 4)

OpenStudy (anonymous):

so the roots are 2x+15 and x-4

OpenStudy (anonymous):

2x + 15 = 0 so one root is x = -15/2 x -4 = 0 so other root is x=4

OpenStudy (anonymous):

oh ok thank you